OpenStudy (anonymous):

Find the real number solutions: 3x^4-27x^2+9x=x^3

OpenStudy (anonymous):

3x^4-x^3-27^2+9x=0 -x^3(-3x+1)9x(-3x+1)=0 (-x^3+9x)(-3x+1)=0

OpenStudy (anonymous):

-3x+1=0 x=1/3 . . .right? And how can i solve for -x^3+9x=0?

OpenStudy (turingtest):

\[3x^4-27x^2+9x=x^3\]\[3x^4-x^3-27x^2+9x=0\]\[x(3x^3-x^2-27x+9)=0\]is what I have so far, I'll continue but if you are right up to -x^3+9x=0 -x(x^2-9)=0 --> x={-3,0,3}

OpenStudy (anonymous):

Thats perfect, thank you so much! :D No need to continue. (:

OpenStudy (turingtest):

cool, I was getting annoyed with this one LOL